Express the series 32 – 20 33 - 21 3k+1 – 2k-1 + 51 52 using the sigma notation.
Answer #1

Clearly, the k th term of the given series is

\frac{3^{k+1}-2^{k-1}}{5^{k}}for\ k=1,2,3,...

Therefore, the given series can be expressed using sigma notation as

\frac{3^{2}-2^{0}}{5^{1}}+\frac{3^{3}-2^{1}}{5^{2}}+...+\frac{3^{k+1}-2^{k-1}}{5^{k}}+...=\sum_{k=1}^{\infty}\frac{3^{k+1}-2^{k-1}}{5^{k}}